OpenStudy (anonymous):

If the height of the building is 250 feet, what is the distance from the top of a building to the tip of its shadow? A. 75.4 feet B. 113.1 feet C. 0.003 feet D. 331.3 feet

OpenStudy (anonymous):

the building

OpenStudy (anonymous):

How would I figure this out?

OpenStudy (anonymous):

That distance, which is the hypotenuse here, we can call "x". sin 49 = 250/x -> x = 250 / (sin 49) So, just use your calculator to get sin 49 and you're all set!

OpenStudy (anonymous):

writing it down :D one second

OpenStudy (anonymous):

Alright I got -262

OpenStudy (anonymous):

Did I do it right?

OpenStudy (anonymous):

oops my mistake i used the bad calculator

OpenStudy (anonymous):

i got 331

OpenStudy (anonymous):

sin 49 approx. = 0.75470958022277199794298421956102 x = 250 / (sin 49) -> x = 250 / 0.75470958022277199794298421956102 x = 331.25324833720283773636185800516

OpenStudy (anonymous):

Good job! You got it! Just saw your post. @Joshuam1

OpenStudy (anonymous):

Sweet :D there are more so I hope you have time

OpenStudy (anonymous):

You're always welcome! As before, new post, so others can join.